1) How many mg of hcl are in 20/ml of the juice with ph 2 molar mass of hcl is 36.5 what is OH minus ions concentration in a solution with PH 2 ?


2) what is physiological range of blood pH ? How the increase of carbon dioxide can influence pH of the blood ? What should be the ratio between the components of the bicarbonate buffer to keep pH of the blood in physiological range ?


3) 0.9% (m/v) Nacl ( 58.5G/mole ) solution is physiological solution . Explain in words or calculation if 10 %of (m/v) saccharose (sucrose) (342g/m) solution can be treated as physiological solution


4 ) There are two water solutions 2 moles/dm*3 glucose (180g/mole ) and 2 moles /dm*3 kcal 74.5G/mole a) explain if the solution is isotonic b) which has lower melting melting point




1
Expert's answer
2021-02-16T07:49:09-0500

1.

pH = - log1 0[ H+]

2 = - log10 [H+]

H+ = 10-2

= 0.01

OH- = 14 - 0.01 = 13.99


2. To maintain homeostasis, the human body employs many physiological adaptations. One of these is maintaining an acid-base balance. In the absence of pathological states, the pH of the human body ranges between 7.35 to 7.45, with the average at 7.40.


So CO2 in the bloodstream lowers the blood pH. When CO2 levels become excessive, a condition known as acidosis occurs. This is defined as the pH of the blood becoming less than 7.35. The body maintains the balance mainly by using bicarbonate ions in the blood.


Ideally, the pH of the blood should be maintained at 7.4. ... Fortunately, we have buffers in the blood to protect against large changes in pH. ... the body) H+ ions and other components of the pH buffers that build up in excess.

I

3.Concentrations of solutions are sometimes given as "percentages." This refers to the mass of solute, in grams, dissolved in a solvent to a volume of 100 ml. "Normal saline" is an aqueous solution of 0.9% NaCl. This means that normal saline can be prepared by measuring out 0.9 g of NaCl and diluting this amount of NaCl to a final volume of 100 ml's in water. This would be the same as diluting 9 g of NaCl to a final volume of 1 liter in water. Normal saline is also known as "physiological saline." Normal saline (0.9% NaCl) is the basis (starting fluid) for intravenous fluid therapy. Normal saline can be directly administered to the bloodstream as it has the same osmolarity as plasma.


4. They are not isotonic because they have different osmotic pressure due to their weight.


The second solution